Circular No. 8854

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V5558 SAGITTARII R. W. Russell, D. K. Lynch, and R. Pearson, The Aerospace Corpora-

tion; M. L. Sitko, University of Cincinnati and Space Science Institute; H. B. Hammel, Space Science Institute; L. Beerman, University of Cincinnnati; and W. Golisch, Infrared Telescope Facility (IRTF), NASA, report on observations made of V5558 Sgr (cf. IAUC 8832) using the IRTF (+ BASS) on July 9 UT. The spectrum was smooth and featureless, roughly following a 2000-K blackbody curve, although there were minor departures from a true blackbody. If the object is a classical nova, then the F-type spectrum (expected near peak brightness) is probably significantly reddened. The narrowband [N] magnitude was ≈ 2.0.

Visual magnitude estimates, most forwarded by E. Waagen, AAVSO: Apr. 21.136 UT, 9.9 (P. Schmeer, Bischmisheim, Germany); 30.090, 9.6 (F. Vohla, Altenburg, Germany); May 5.347, 9.3 (L. Shotter, Uniontown, PA, U.S.A.); 18.504, 9.0 (S. Kerr, Glenlee, Qld., Australia); 24.072, 8.6 (A. Amorim, Florianapoulos, Brazil); June 5.944, 8.5 (A. Kammerer, Malsch, Germany); 11.987, 8.4 (J. Speil, Walbrzych, Poland); 24.122, 8.4 (J. Bortle, Stormville, NY, U.S.A.); 26.861, 8.1 (S. Bolzoni, Busto Arsizio, Italy); 29.396, 7.5 (H. Matsuyama, Kanimbla, Qld., Australia); July 3.093, 7.5 (Bortle); 7.903, 6.7 (T. Scarmato, Calabria, Italy); 11.866, 6.6 (K. Hornoch, Vohancice, Czech Republic).

THE EDGAR WILSON AWARD 2007 The Smithsonian Astrophysical Observatory announces that the 2007

Edgar Wilson Award for the discovery of comets (cf. IAUC 6936, 8730) is being divided among the following three individuals for four different comets: John Broughton (Reedy Creek, Qld., Australia) for C/2006 OF_{2} (cf. IAUC 8756); David H. Levy (Tucson, AZ, U.S.A.) for P/2006 T1 (cf. IAUC 8757); and Terry Lovejoy (Thornlands, Qld., Australia) for C/2007 E2 (cf. IAUC 8819) and C/2007 K5 (cf. IAUC 8840).

COMET 188P/LINEAR-MUELLER Comet P/1998 S1 = P/2007 J7 (cf. IAUC 8853) has been assigned the

permanent number 188P (cf. MPC 60060).

2007 July 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green